Please help me solve this equation..solve where u is the real number(u-9)^2-27=0
Answer by stanbon(75887)

Please help me solve this equation..solve where u is the real number(u-9)^2-27=0
---------------
(u-9)^2 = 27
-----------
u-9 = +3sqrt(3) or u-9 = -3sqrt(3)
u = 9 + 3sqrt(3) or u = 9 - 3sqrt(3)
=========================================
Cheers,
Stan H.